POLYMER CHEMISTRY For polystyrene with Mn = 50,000g/mol: estimate end-to-end distance for ideal fully extended(it is a contour length) and coiled conformation assuming that the macromolecules behave as a freely rotating,freely jointed,volumeless chains.
Answer #1

M=50,000 g/mol.

The repeating unit, C8H8, has a molar mass of 104.15 g/mol.

Each unit has two backbone bonds. So the molar mass per backbone bond=104.15/2=52.075 g/mol

So the number of bonds, n=50,000/52.075=960.

The bond-length (C-C bond), l ~ 1.54 A0.

Therefore the mean-square end-to-end distance <R2>=C n l2=9.5*960*1.542 (A0)2=21629 (A0)2

(C=9.5 is taken for polystyrene)

So the mean end-to-end distance of a fully extended chain is given by

<R2>1/2=147 A0.